You are an assistant to the CFO and have collected the following data to conduct the analysis.

  • The company is subjected to a marginal tax rate of 35%.
  • The company can issue a 20-year, 7.6% semi-annual coupon bond at $1,219. New bonds would be privately placed with no floatation cost.
  • The company’s preferred stock currently sells for $30 per share. It pays a fixed dividend of $1.8 per share.
  • The company’s common stock currently sells for $40 per share. The most recently paid dividend was $1 per share. Dividends are expected to grow at a constant rate of 5% in the foreseeable future.
  • A flotation cost of 10% would be required to issue new common stock.
  • The company’s stock beta is 1.875, the market risk premium is 3% and the risk-free rate is 2%.
  • The company’s capital structure consists of 30% debt, 5% preferred stock, and 65% common equity.

(1)   Estimate the cost of the following capital components

  • Cost of Debt
  • Cost of Preferred Stock
  • Cost of Retained Earnings—based on DCF approach
  • Cost of Retained Earnings—based on CAPM approach
  • Cost of New Common Stock ---based on DCF approach

(2)   Calculate the firm’s WACC, assuming it does not want to issue new common stock.

(3)   Calculate the firm’s WACC, assuming the firm expands so rapidly that it must issue new common stock.

"not using excel"
Answer #1

Following information is given in the question –

  1. Semi-annual coupon rate of bond is 7.6%
  2. Preference shares current price is $30 per share
  3. Preference shares current price is $30 per share
  4. Preference shares current dividend is $1.80 per share fixed
  5. Common equity shares current price is $40 per share
  6. Common equity shares current dividend is $1 per share
  7. Growth rate of divided is 5%
  8. Flotation cost to issue new common stock is 10%
  9. Company’s stock beta is 1.875, the market risk premium is 3% and the risk-free rate is 2%
  10. Company’s capital structure consists of 30% debt, 5% preferred stock, and 65% common equity
  11. Marginal tax rate is of 35%

Answer – 1

Cost of Debt (Kd) = After tax Interest on Bonds

= 7.6 * 2 (1 – Tax Rate)

= 7.6 * 2 (1 – 0.35)

Kd = 9.88%

Cost of Preference shares (Kp) = (Current Dividend per share / Current Market Price per share) * 100

= (1.80 / 30) * 100

= 6%

Cost of Retained Earnings (Ke)(using DCF approach) = Cost of Common Equity

DCF approach can be used by using the Dividend Growth Model which states as below :

P = D1 / (Ke – g)

Where –

P is Common equity shares current price = $40 per share

D1 is Common equity shares dividend after one year

                D1 = Current dividend + one year growth

                      = $1 (1 + 0.05)

                      = $1.05 per share

Ke is Cost of Common Equity

g is growth rate = 5% or 0.05

On putting these figures in the formula, we get –

P = D1 / (Ke – g)

40 = 1.05 / (Ke – 0.05)

On cross multiplying

40 (Ke – 0.05) = 1.05

40 Ke – 2 = 1.05

40 Ke = 1.05 + 2

40 Ke = 3.05

Ke = 3.05 / 40

Ke = 0.07625 or 7.625%                

Hence,

Cost of Retained Earnings (using DCF approach) = Cost of Common Equity (Ke)

                                                                                = 7.625%

Cost of Retained Earnings (Ke) (using CAPM approach) = Rf + β (Rm – Rf)

Where –

Rf is the risk-free rate of 2%

β is company’s stock beta of 1.875

Rm = market risk premium + risk-free

       = 3% + 2%

       = 5%

On putting these figures in the formula, we get –

Cost of Retained Earnings (using CAPM approach) = Rf + β (Rm – Rf)

                                                                                   = 2 + 1.875 (5 – 2)

                                                                                   = 2 + 1.875 * 3

                                                                              Ke = 7.625%

Cost of New Common Stock (based on DCF approach) = Cost of common stock + flotation cost

                                                                                          = 7.625 (1 + 0.10)

                                                                                           = 8.3875 %

Answer – 2

Weighted Average Cost of Capital (WACC) (assuming company does not want to issue new common stock) = Kd * W1 + Kp * W2 + Ke * W3

Where –

Kd = 9.88%

Kp = 6%

Ke = 7.625%

W1 = 30% or 0.30

W2 = 5% or 0.05

W3 = 65% or 0.65

On putting these figures in the formula, we get –

WACC = Kd * W1 + Kp * W2 + Ke * W3

            = 9.88 * 0.30 + 6 * 0.05 + 7.625 * 0.65

            = 2.964 + 0.30 + 4.95625

            = 8.22%

Answer – 3

Weighted Average Cost of Capital (WACC) (assuming company want to issue new common stock) = Kd * W1 + Kp * W2 + Ke * W3

Where –

Kd = 9.88%

Kp = 6%

Ke = 8.3875%

W1 = 30% or 0.30

W2 = 5% or 0.05

W3 = 65% or 0.65

On putting these figures in the formula, we get –

WACC = Kd * W1 + Kp * W2 + Ke * W3

            = 9.88 * 0.30 + 6 * 0.05 + 8.3875 * 0.65

            = 2.964 + 0.30 + 5.4519

            = 8.72%